### v1

# Character is Destiny: Can Large Language Models Simulate Persona-Driven Decisions in Role-Playing?

Abstract
Can Large Language Models (LLMs) substitute humans in making important decisions? Recent research has unveiled the potential of LLMs to role-play assigned personas, mimicking their knowledge and linguistic habits. However, imitative decision-making necessitates a more nuanced understanding of personas. In this paper, we benchmark the ability of LLMs in persona-driven decision-making. Specifically, we investigate whether LLMs can predict characters’ decisions provided with the preceding stories in high-quality novels. Leveraging character analyses written by literary experts, we construct a dataset LifeChoice comprising 1,401 character decision points from 395 books. Then, we conduct comprehensive experiments on LifeChoice, with various LLMs and methods for LLM role-playing. The results demonstrate that state-of-the-art LLMs exhibit promising capabilities in this task, yet substantial room for improvement remains. Hence, we further propose the CharMap method, which achieves a 6.01% increase in accuracy via persona-based memory retrieval. We will make our datasets and code publicly available.
Figure 1: An example of NextDecisionPrediction. LLMs generate a profile and memory from the historical data before the decision node in the storyline and then make life choices from the perspective of the character under the current scenario.
1 Introduction
If we can thoroughly understand all the causes, the future is not an unpredictable miracle but an inevitable result of the past.
– Voltaire
Decision-making is a vital part of human activities, and every important decision reflects a person’s history. The recent progress of large language models (LLMs)OpenAI (2023); Touvron et al. (2023); Team (2023) have enabled them to simulate the personas of assigned individuals, inheriting knowledge, linguistic patterns, and even personalities Park et al. (2023); Gao et al. (2024); Wang et al. (2024). However, a key challenge towards persona-playing LLMs lies in persona-driven decision-making, i.e., whether LLMs could make decisions consistent with the intended personas, as depicted by their past behaviors. This capability is indispensable for a broad range of applications, such as understanding human behavior patterns Fellows (2004), evaluating role-playing models Wang et al. (2023, 2024), and constructing personal assistant Xu et al. (2022); Salemi et al. (2024).
Previous work has studied language models in understanding characters in fictional works, e.g., guessing a character’s identities  (Sang et al., 2022) or possible actions (Mostafazadeh et al., 2016) from the context. However, their studies are limited to short and straightforward contexts. In contrast, recent persona-assigned LLMs are built on massive persona data, making their decision-making process much more complicated. Previous analysis in persona-assigned LLMs generally focuses on linguistic styles (Wang et al., 2023), knowledge memorization (Shao et al., 2023) and personality traits (Wang et al., 2024). These studies are restricted in evaluating surface-level behaviors, leaving it unclear whether persona-assigned LLMs can replicate the thought process and mindset for well-considered actions, i.e., their decision-making. Hence, the persona-driven decision-making of LLMs remains an important yet underexplored question.
In this work, we propose NextDecisionPrediction, a novel task to benchmark the ability of LLMs in simulating the decision-making process of the target personas. As shown in 1, each sample in NextDecisionPrediction is a key decision point of a character, with character data preceding the point as the input and their decision as ground truth for the output. In delicately plotted novels, the key character decisions are typically foreshadowed, leading to their destiny. Therefore, we construct LifeChoice, a multi-choice dataset based on literature analysis for fictional characters, contributed by literature experts from the Supersummary [^1] https://www.supersummary.com/. The construction of LifeChoice includes three steps: decision point detection, distractor generation, and motivation extraction. To predict the decision, LLMs need to identify and reason over the persona’ historical data related to the current scenario.
We evaluate persona-assigned LLMs on LifeChoice in various settings. Following prior research (Li et al., 2023; Wang et al., 2023), we first craft character profiles containing descriptions and memories, which compress the preceding character data. Then, we build persona-assigned LLMs based on character profiles. With extensive experiments, we find that existing persona-assigned LLMs achieve an accuracy up to 76.95% on LifeChoice, and both well-summarized character descriptions and accurate memory retrieval are important to persona-assigned LLMs. Furthermore, we propose CharMap, which adopts persona-based memory retrieval and gains a 6.01% accuracy improvement over the baselines.
In summary, our contributions include:
  • We propose NextDecisionPrediction, the first task to evaluate the decision-making capabilities of persona-assigned LLMs.
  • We construct LifeChoice, a dataset for NextDecisionPrediction. It is constructed from life choices of characters in literature work, sourced from literature analysis from experts. Besides, We propose CharMap, which adopts persona-based memory retrieval to improve LLM role-playing.
  • With extensive experiments, we find that state-of-the-art persona-assigned LLMs exhibit promising capabilities in persona-driven decision-making. Additionally, we analyze methods and future directions for further improvement.
2 Related Work
Character Role-Playing
Early research on character-related studies focuses on character understanding. Brahman et al. (2021) attempts to predict a specific character through the text of the novel. Yu et al. (2022) provides dialogues from movie scripts for the model to examine and then asks it to identify the character who speaks each passage. With the enhancement of model abilities, some work attempts to make the model simulate complex role-playing. Li et al. (2023) analyzes 32 anime characters using 54k dialogues and personality traits. They use sentence embeddings for dialogue selection and evaluation. Zhou et al. (2023) uses identity, interests, and relationships, collecting AI behaviors for imitation and using character data for fine-tuning. They evaluate model consistency and linguistic style. Wang et al. (2023) creates a dataset for script characters and evaluates role-playing quality based on speaking style imitation and role-specific knowledge. These studies make a chatbot for a certain character, but they focus more on imitating the character from the perspective of dialogue, which is a shallow imitation. We aim to role-play from the perspective of behavior and decision-making. This form tests the model’s understanding of the role more.
Personal LLM assistants
With the rapid development of artificial intelligence technology, there are now many personal intelligent agents embedded in mobile devices, providing personalized services through analyzing user data and equipment Kaplan and Haenlein (2019); Hoy (2018). These agents can model the user’s profile and preferences through the user’s historical data Gurrin et al. (2014); Dodge and Kitchin (2007), such as extracting personality from the user’s record text Majumder et al. (2017); Štajner and Yenikent (2020), reading emotions from the user’s image data Jaiswal et al. (2020); Zad et al. (2021), modeling preferences from historical interaction information Tang et al. (2019); Li et al. (2018), and pushing notifications from smart phones Li et al. (2018). These memories can enhance the model’s decision-making and reasoning, bringing a better personal experience for users. However, obtaining real user memory data is difficult and violates privacy. We model characters from historical data in high-quality novel texts, allowing the model to restore the real choices in the storyline based on the previous text, providing the first benchmark for the wide testing of personal intelligent agents.
3 Datsets and Task Setups
3.1 Dataset Construction
Book: Les Misérables
Character: Jean Valjean
Input:
In 1815 Monsieur Charles-François-Bienvenu Myriel was Bishop of Digne. He was then……Jean Valjean had retaken his seat at the end of the table. He had resumed his thinking. His reflections gave him a sort of frightening aspect. He was subject to one of those violent inner tearings, which was not unknown to him.
Scenario:
In the courtroom, an innocent man was wrongfully accused of being him because he bore a resemblance to Jean Valjean. If Jean Valjean did not come forward, this innocent man would be sent to the gallows in his place. At this time, Jean Valjean had transformed his identity and become a respected town mayor, and he had also adopted a young girl named Cosette, with whom he had a new life.
Question:
You will play the role of Jean Valjean. What will you choose to do when you discover that man is about to be convicted due to being mistaken for you?
Options:
A. Keep silent, letting an innocent person take the punishment in one’s place.
B. Persuade the person to run away, in order to protect both from the disaster of jail.
C. Go to court and reveal the truth, sacrificing oneself to save the innocent person.
D. Look for legal loopholes, trying to save both the person and oneself.
Correct Answer: C
Motivation:
[Values and Beliefs] Jean Valjean is a person who values honesty and justice, possessing a strong sense of morality and righteousness. He decides to turn himself in to save another innocent person, fulfilling his inner need for morality and justice.
Table 1: Case study of LifeChoice. A complete set of data includes book, character, scenario, question, options, correct answer, motivation, and input.
Dataset Source Context Length Task Format Has Explanation
TVSHOWGUESS TV show transcripts ~50k Character Identification
ROCStories Commonsense short stories ~100 Character Behavior Prediction
LiSCU Literature ~1000 Character Identification
LIFECHOICE Literature ~150k Character Behavior Prediction
Table 2: Comparison between LifeChoice and previous character understanding benchmarks: data source, context length, task format, and whether the benchmark has explanations.
We build our dataset on the Supersummary website, which provides high-quality summaries and analyses of novels by literary experts. We connect with this website and utilize their full book summaries and character analyses. With these resources, we design life-choice questions for key characters using GPT-4. The dataset construction comprises the following three main steps:
Selecting Available Novels
To prevent data leakage, we filter all novels on the site using the following criteria: (1) The narrative must exclude non-fiction genres, such as biographies or documentary literature. (2) The narrative perspective must be in the first or third person. (3) The progression of narrative time should be linear, avoiding stories with complex timelines or flashbacks. For each book that passes these filters, we apply GPT-4 to parse the character analysis provided by literary experts, extracting important decision nodes and their motivations. If these motivations relate to events from earlier chapters, we identify the original chapter text through comprehensive summaries.
Constructing Multiple-Choice QA Data
We input the full novel summaries, characters’ decision nodes, and potentially influential chapter texts into GPT-4. Our goal is to generate multiple-choice questions that capture the complexity of the characters’ decision-making processes. Every option is designed to seem plausible, while only one choice—the one made in the original book—is correct.
Manual Filtering
We invite ten native English-speaking university students to filter all the data. We supply the annotators with each data sample’s corresponding summaries and human analyses, enabling them to determine whether the multiple-choice questions created by the model are challenging and reasonable. The detailed annotation rules are available in Appendix B.1.
In the end, we collect 1,401 characters from 396 books and their corresponding life choices. Table 1 shows a complete data example.
Figure 2: Statistics of motivation types in LifeChoice, with the first words for each motivation type.
3.2 Dataset Analysis
We refer to the drama theory of Aristophanes Sommerstein (2013); Silk (2002) as the system prompt, use GPT-4 to classify the motivations for character decisions into two meta-motivations and several accompanying sub-motivations:
Character-driven motivation
Character-driven behavior revolves around the character’s inner world, personality, and transformation. Sub-motivations of character-driven behavior include Personality and Traits, Emotions and Psychological State, Social Relationships, Values and Beliefs, and Desires and Goals.
Plot-driven motivation
Plot-driven behavior stems from a series of external events and the unfolding of conflicts. Characters often react passively within a larger narrative structure, with their actions led by external events. Sub-motivations of plot-driven behavior include External Conflicts, Tasks and Goals, Puzzles and Secrets, Pursuits and Escapes, Exploration and Discovery, Power and Control, and Intrigue and Betrayal.
Note that each topic will only be assigned one category of motivation. Figure 2 shows the proportion of different motivations. Detailed introductions for each sub-motivation are in Appendix A.2.
3.3 Task setups
Given an input $\mathcal{X}=(\mathcal{D},\mathcal{S},\mathcal{C},\mathcal{Q},\mathcal{A})$, comprising the preceding textual content from the original book before the decision node $\mathcal{D}$, the current scenario $\mathcal{S}$, the focused character $\mathcal{C}$, a multiple-choice question $\mathcal{Q}$ outlining a decision faced by that character, and a set of candidate answers $A=\{A_{i}\}_{i=1}^{4}$. The ultimate goal is to identify the correct choice $\mathcal{Y}$ that aligns with the character’s decision in the narrative.
This task can be formulated as $P(\mathcal{Y}|\mathcal{X})$. For evaluation, we directly use the accuracy of multiple-choice QA. As shown in Table 2, compared to other character understanding tasks, the LifeChoice requires understanding the characters through an excessively long context to make selections. Similar to Personal LLM assistants, the model needs to locate relevant information in a vast amount of sparse personal data related to the current scene. This behavior demands a more profound understanding of the characters.
4 Experiments
Profile Construction Reasoning Model Task QA Profile Score
ACC +motivation
Only description
Model DescriptionWu et al. (2021) gpt-4-1106-preview 48.43 89.24 1.84
Human Description gpt-4-1106-preview 60.02 95.47 2.23
Only memory
Model Retrieval(BM25) gpt-4-1106-preview 26.08 86.88 1.07
Model Retrieval(Embedding) gpt-4-1106-preview 38.66 89.24 1.75
Description+Memory
Model Description+Model Retrieval gpt-3.5-turbo-1106 51.62 96.39 1.96
gpt-4-1106-preview 58.92 93.46 2.05
[…]
Claude2.1 68.02 95.24 2.42
gpt-3.5-turbo-1106 67.02 98.85 2.35
gpt-4-1106-preview 70.94 100.00 2.47
[…]
gpt-4-1106-preview 76.95 100.00 2.67
Table 3: Results of different LLMs performance on NextDecisionPrediction. ACC refers to the decision accuracy. +motivation refers to the results after adding manually written motivation by experts to the constructed profile, and Profile Score refers to the score given by GPT-4 for whether the profile contains the meaning of motivation.
Since our inputs generally exceed 100k, it is difficult for LLMs to handle them directly. Therefore, our method is divided into two steps: First, we compress the input to under 5k tokens(the maximum token count for human description) to construct a character profile that complies with most LLMs’ token restrictions. Second, we feed the constructed profile and the question into different LLMs for responses.
4.1 Character Profile Construction
For the first step, we refer to the role-playing model Li et al. (2023); Zhou et al. (2023) and propose two baseline methods. The first method summarizes all the context into a description of the corresponding character, composed of the character’s basic situation and storyline. This provides a holistic portrayal of the character during role-playing. The second method focuses on retrieving memories of the current scene, which could offer more detail. Below, we detail the methods we use:
Description Construction
We adopt the method by Wu et al. (2021) to summarize the characters’ profiles. Considering the model’s capabilities and costs, we use GPT-3.5 to replace GPT-3 in the original paper. We divide $\mathcal{D}$ into chunks, summarize them separately, and then merge them. After several iterations, we generate the final summary. We also use the manually written character storylines in Supersummary as the human description.
Memory Retrieval
We use two retrieval methods to find memory from $\mathcal{D}$ relevant to the current scenario $\mathcal{S}$. The first is the BM25 algorithm Robertson et al. (2009), and the second uses embedding-based retrieval. For the embedding retrieval, we break $\mathcal{D}$ into chunks of 1k tokens and match them with vectors extracted by the embedding model for $f(\mathcal{S},\mathcal{Q})$. Both retrieval results were kept within 5k tokens. For the embedding model, we use OpenAI’s text-embedding-ada-002Neelakantan et al. (2022) model.
Description & Memory
Using description and memory alone may result in missing information thus we attempt to concatenate the results of both methods. For the merged text, we keep it within 5k tokens.
4.2 Reasoning the Answer
After compressing the original input $\mathcal{X}$ into a Character Profile, we feed it into the LLMs. For the individual methods, we respectively use GPT-4 as the final reasoning model. For the joint method, we experiment with the following LLMs: Mixtral-8x7B-MoE Jiang et al. (2024), Claude2.1 Anthropic (2023), GPT-3.5-Turbo OpenAI (2022), GPT-4-Turbo OpenAI (2023). For all these models, we adopt the official instruction formats where available.
5 Analysis
Raw text Model desc. Human desc.
+Model retr. +Model retr.
GPT-4 - 58.92 70.94
human 92.01 65.82 72.78
Table 4: Results of the human evaluation. The inputs provided to humans are the raw text, model description + model retrievals, and human description + model retrievals, respectively.
Restoration Level Raw input Entity replacement
completely consistent 27.5% 0%
partially consistent 47.5% 37.5%
entirely different 25.0% 62.5%
Table 5: Analysis of Data Leakage Mitigation through Entity Replacement. Three restoration levels represent the model’s ability to generate the original plot before and after replacement, and the percentages represent the proportions of different restoration levels.
In the experiments, we wish to answer three research questions: RQ1) Can LLMs make decisions based on historical data? RQ2) What influences the decision-making of LLMs? RQ3) How can the decision-making of LLMs be improved?
5.1 Can LLMs make decisions based on historical data?
How do models make decisions based on historical data?
Table 3 reports the accuracy results of baseline methods on the future prediction task in our dataset. Several observations can be made: First, the performance of the combined methods outperforms the use of a single one, suggesting that both holistic and detailed data of characters are important in final decision-making. Second, when provided with a gold explanation, the accuracy consistently exceeds 90%, indicating the rationality of these explanations in the data. Third, the performance gap among different LLMs is not significant while reasoning the answer. This indicates that the main factor for the result is the generated profile, rather than reasoning ability.
How do humans make decisions based on historical data?
We invite three native English-speaking university students to participate in a test where we select six novels they have never heard of before. Each novel has between 5 to 10 characters and their corresponding multiple-choice questions. We test three groups, allowing them to make choices after reading all content before the decision nodes: model-generated description + embedding-retrieved memories or human description + embedding-retrieved memories, as shown in Table 4. The results indicate that humans have a strong ability to make choices after reading the original text. When provided with the compressed content, human performance is slightly better than GPT-4. This indicates a considerable scope for enhancing our method, particularly in acquiring information from the original text that impacts the characters.
Do LLMs know the plot following the node in advance?
Data leakage poses a significant challenge for our benchmarks since the novel data we utilize might have appeared in the pre-training set. We aim for the model to forget relevant plots as much as possible. To this end, we adopt the entity replacement strategy. During the evaluation, we replace character names with placeholders like [character A]. Further, we replace placenames and some distinctive phrases that could lead to data leakage, such as the iconic sentence “For you, a thousand times over” from the novel “The Kite Runner”. We utilize GPT-4 to identify these potential leakages and have applied rules to replace them.
To test the effectiveness of this approach, we input the human summary before and after entity replacement, and ask the model to output the subsequent content, evaluating its capability to restore the plot from the book. The levels of restoration are categorized as “completely consistent”, “partially consistent”, and “entirely different.” We conduct a manual evaluation of 40 characters from 20 books. Specific evaluation details can be found in Appendix B.2. As shown in Table 5, After replacing the entities, the model no longer produces content that is “completely consistent”, and the overall situation of data leakage is significantly alleviated.
5.2 What influences the decision-making of LLMs?
Figure 3: Heatmap of the impact of motivation types on the results. The results are predicted from the model-generated description, the model-retrieved memory, the combination of both, and CharMap. The reasoning model uses GPT-4.
Does the type of motivation affect the model’s decision-making?
In line with the motivation types presented in Section 3.2, we examine how different types of motivation influence characters’ decision-making. For profiles, we evaluate three scenarios: using only model-generated descriptions, embedding-retrieved memories, and a combination of both. For reasoning, we use GPT-4 uniformly. The results are shown in Figure 3. We find that for all methods, tasks requiring coherent reasoning, such as puzzles and mysteries, are not well answered. This might be because these questions need multi-step reasoning and details from various memories. Moreover, plot-driven questions have lower accuracy when only descriptions are used for the profile. Conversely, character-driven questions are challenging to answer when relying only on memories. We believe this is because character summaries in descriptions better capture the overall essence of the characters, while memories provide direct access to relevant events.
Figure 4: The result of the impact of different novel genres on accuracy.
Does the genre of the novel affect the model’s decision-making?
We use the genre tags from novels on the website to analyze the accuracy of character selection across different genres. We conduct experiments on the combination of human description + embedding-retrieved memories, and the reasoning model using GPT-4. As depicted in Figure 4, the accuracy of science fiction, fantasy novels, and romance novels is quite high. This could be due to the characters in these novels often being stylized or having fixed creative patterns and archetypes. In contrast, crime and mystery novels perform poorly, which might be because they involve complex logical chains, and characters in these novels frequently take abnormal actions. Further details about each genre and the complete table can be found in Appendix A.1.
Are expert-annotated motivations considered gold interpretable motivations?
We need to ensure that the problems and motivations are corresponding. As shown in Table 3, when we use these motivations as part of the input, the accuracy of the QA task exceeds 85%. This demonstrates that the motivations effectively aid the model in making selections. In addition, we adopt the evaluation method used in recent research Liu et al. (2023); Li et al. (2024), utilizing GPT-4 to evaluate whether the model-generated profiles contain human motivations. We assign scores from 1 to 3 for the model in this regard:
Score 1: The paragraph fails to convey the core information of the target sentence, exhibiting significant misunderstandings or omissions.
Score 2: The paragraph conveys the main information of the target sentence but misses some details or has minor inaccuracies.
Score 3: The paragraph completely and accurately conveys all the key points of the target sentence. We calculate average scores using different methods.
As Table 3 shows, a higher score suggests that the model-generated profile more accurately represents the information from the manually created motivations. Moreover, higher scores correspond to increased final accuracy, which reiterates the validity of the manual motivations. To uphold the quality of the GPT-4 evaluation, we select a sample of data for human evaluation and comparison and compute the Pearson Correlation Coefficient Cohen et al. (2009) across both evaluation types. The value reaches 0.81, indicating that GPT-4’s evaluation capabilities on this task are on par with humans. The specific evaluation process can be found in Appendix B.3.
Figure 5: Analysis of whether character selection will change. The x-axis represents the input length relative to the node truncation.
Do characters at different times make the same choices?
If faced with the decisions of years past at this moment, would you make the same choices? We conduct a study on this matter. Specifically, we randomly sample 40 characters, half character-driven and half plot-driven. We split the content preceding the decision points into five equal sections and used these various content lengths as input. We conduct experiments on the combination of human description + embedding-retrieved memories, and the reasoning model uses GPT-4. As shown in Figure 5, in the early stages, the accuracy of most characters’ decisions is close to random (25%), potentially due to insufficient information. As more information becomes available, the characters’ decisions tend to be closer to the correct choice. For character-driven decisions, accuracy tends to be stable. For plot-driven, the accuracy rate may change abruptly. This could be due to the relatively stable characteristics of a character, while some sudden events may greatly influence the final choices of the character.
Figure 6: An overview of CharMap, a two-step scenario-specific character profile building approach. It locates content in memory that is more consistent with the current scenario through the connection between description and question.
5.3 How can the decision-making of LLMs be improved?
Based on the conclusion in section 5.2, the model often performs poorly in answering questions requiring multiple details or multi-hop reasoning, possibly due to the overly dispersed memory retrieved. We propose the CHARacter MAPping Profile Synthesis (CharMap), constructing a more scenario-specific profile in two steps. As shown in Figure 6, first, after obtaining the description generated by the model, we input it along with the question into the model, asking the model to locate the plot in the description relevant to the current scene based on the question. Second, we use these episodes as queries to retrieve related memory, and then input them into the inference model together with the description. This can take advantage of the overall description of the character storyline in the description, thereby better retrieving related memories. As shown in Table 3, the accuracy after using CharMap is 6.01% higher than directly concatenating the description with memory. As seen from Figure 3, the accuracy rate of each category of questions improves after using the CharMap, especially the Puzzles category that requires multi-hop reasoning.
6 Conclusion
In this work, we propose NextDecisionPrediction, the first task to evaluate the persona-driven decision-making of LLMs, testing whether LLMs can reconstruct actual storylines through historical character data. We construct a dataset LifeChoice, including 1,401 characters from 396 books and their significant decisions. We analyze the performance of LLMs on this task from various aspects, such as novel genres, character motivations, and model solution approaches. Based on this, we propose a method, CharMap, which achieves better results by constructing a character profile more relevant to the current scene in two steps.
Decision-making is an important and complex behavior of humans, and we hope to explore the possibility of LLMs making personal decisions. This could establish important evaluation criteria for role-playing models and Personal LLM assistants.
Limitations
The partial evaluation method we proposed is dependent on GPT-4, which could be biased towards GPT-4 generations. Furthermore, although we have alleviated the data leakage problem through entity replacement, there is still a small amount of data that may have data leakage, especially in some classic novels. Finally, our dataset is constructed through the decision of high-quality novel characters. However, compared to human choice, this part of the data is not sparse or challenging enough. We hope to construct real human decision-making data while ensuring privacy.
Appendix A Dateset Details
A.1 Categories of novel
Below is a complete classification of novel genres, from the literary experts at the Supersummary website:
Mystery Novels: The mystery genre includes general mystery, noir mystery, historical mystery, police procedural mystery, and supernatural mystery.
Thriller Novels: The thriller genre includes supernatural thrillers, historical thrillers, environmental thrillers, medical thrillers, legal thrillers, political thrillers, military thrillers, and espionage stories.
Science Fiction Novels: Science fiction stories take place in the future or the past but are almost always set in a dimension different from our present. They are characterized by entirely new, imagined realities and universes, where the setting is indispensable. High technology also plays an important role in these stories. Space opera, romantic science fiction, military science fiction, alternate history, dystopian and utopian tales, as well as steampunk, are considered sub-genres of science fiction.
Romance Novels: Romance novels feature romantic relationships between at least two people, characterized by tension and desire. Romance novel themes include supernatural romance, contemporary romance, historical romance, western romance, gothic romance, regency romance, and romantic suspense.
Fantasy Novels: Fantasy stories are centered around mythical kingdoms and magic. Fantasy novel genres include contemporary fantasy, traditional fantasy, horror fantasy, weird fantasy, epic fantasy, historical fantasy, dark fantasy, urban fantasy, and anime fantasy.
Action Adventure Novels: Action-adventure novels place the protagonist in various realistic dangers. This is a fast-paced genre where the climax should provide some form of thrill for the audience or reader.
Speculative Novels: Speculative fiction is characterized by overlapping with our world but differing in key aspects, introducing "what if" scenarios.
Mystery Thriller Novels: Mystery thriller stories are usually filled with suspense, with one or more characters’ lives in danger. In gripping scenes, these characters are often chased and manage to escape narrowly.
Young Adult Novels: Young Adult fiction, commonly abbreviated as YA, is intended for teenagers aged 12-18. Most YA novels feature coming-of-age stories, often with elements of science fiction or fantasy.
New Adult Novels: New Adult novels target college-aged adults and usually explore stories of first adventures on one’s own.
Horror and Supernatural Novels: Horror, supernatural, and ghost story genres aim to scare the reader and audience by playing on common fears. The protagonist usually has to overcome supernatural threats, and the stories often include supernatural elements.
Crime Mystery Novels: Crime mystery stories focus on a central problem or crime to be solved, or a mysterious event that must be answered. Throughout the story, the reader or audience and characters are given clues that help the protagonist eventually find the solution.
Detective Novels: In detective fiction, a common element is a police officer or detective embarking on solving a crime. The plot is filled with evidence gathering, forensic studies, and legal drama.
Historical Novels: Historical novels are fictional stories set against the backdrop of real historical events or historical settings. Historical fiction may also portray real historical figures.
Western Novels: Stories with a western theme take place in the old times of the American West, filled with adventure, cowboys, and pioneers. There are also Italian western novels, Asian western novels, space westerns, and other stories about the American West.
Family Saga Novels: Family saga novels typically tell the stories of several generations of family members dealing with family affairs, family curses, and family adventures. These stories usually follow a timeline and deal with conflicts in the present.
Women’s Novels: Women’s fiction plotlines revolve around the challenges and crises that women face in real life, including interpersonal relationships, work, family, politics, and religion.
Magical Realism Novels: Magical realism stories take place in the real world but have characters who take magical elements for granted. These magical elements do not exist in real life, but they are perfectly normal in the realm of magical realism.
A.2 Categories of motivations
Below are the motivations for each topic and their corresponding proportions:
Character-driven motivation
Character-driven narrative is centered on the inner world, growth, and transformation of characters. In character-driven stories, the progression of the plot and the resolution of conflicts are often propelled by the characters’ personalities, desires, fears, and psychological development. Such stories typically delve deeply into the characters’ mental states and development, focusing on how characters influence each other and how their actions reflect their inner emotions and thoughts. The choices and changes of the characters serve as the main engine for the story’s development, influencing the direction of the plot. Sub-motivations of character-driven behavior include:
Personality and Traits: (27.12%) These refer to a character’s characteristics such as being introverted, extroverted, brave, or guilt-ridden, which influence their choices and lifestyle.
Emotions and Psychological State: (7.53%) A character’s emotional responses, psychological traumas, or sense of personal well-being are key elements that drive the story forward.
Social Relationships: (6.31%) The character’s status and changes in family, love, friendship, or other social connections can propel the story’s development.
Values and Beliefs: (27.12%) The character’s moral convictions, religious beliefs, or life philosophy can serve as motivation for action.
Desires and Goals: (7.22%) Personal desires, career aspirations, or specific life goals of a character are pivotal in advancing the plot.
Plot-driven motivation
Plot-driven narrative emphasizes the creation and resolution of external conflicts in the story. In such stories, the driving force of the plot comes from a series of events and conflicts themselves, while characters are often the responders to these events. Plot-driven stories typically highlight tense drama, complex plot structure, and frequent changes in external actions, rather than changes in the character’s internal world. In this type of narrative, characters may act in response to the demands of the plot, rather than the plot following the development of the characters’ inner world. Sub-motivations of plot-driven behavior include:
External Conflicts: (8.76%) Conflicts from the outside world, such as war, natural disasters, or social upheaval, can propel the plot.
Tasks and Goals: (4.7%) Tasks or specific goals that characters must accomplish often become the driving force behind the story’s progression.
Puzzles and Secrets: (7.22%) Secrets that need revealing or mysteries that need solving can form the core of a story.
Pursuits and Escapes: (4.25%) Characters might chase something (e.g., power, wealth, knowledge) while avoiding or fleeing from certain situations (e.g., pursuit, personal past).
Exploration and Discovery: (3.66%) Characters’ adventures or discoveries in new realms (physical, scientific, or spiritual) can move the plot forward.
Power and Control: (4.81%) The pursuit or struggle for power and control often serves as motivation for characters.
Intrigue and Betrayal: (4.09%) Complex plots and betrayals can catalyze the progression of the story.
Appendix B Manual Annotation
For all individuals involved in the annotation, we provide compensation based on the local minimum hourly wage.
B.1 Manual Filtering Rules
This is a supplement to Section 3.1. After constructing the multiple-choice question data using GPT-4, we perform manual filtering.
For each annotator, we provide novel summaries and character analyses written by human literature experts on the Supersummary website. Each annotator is asked to score the questions constructed by GPT-4 based on the following evaluation criteria:
1. Comprehensiveness
Rule 1.1: Evaluators must ensure that each multiple-choice question fully considers the character’s background, context, and motivation. The questions should reflect the true decisions and experiences of the character within the narrative.
Scoring Guide:
2 points: The question is detailed and comprehensive, aligning perfectly with the character’s background and motivation.
1 point: The question aligns generally but is missing key aspects of the character’s background information or motivational nuances.
0 points: The question significantly misaligns with the character’s background or motivation.
2. Logical Consistency
Rule 2.1: Evaluators should assess the internal consistency and plausibility of the question within the narrative thread. The content and structure of the multiple-choice question must be consistent with the plot and the character’s logical decision-making process.
Scoring Guide:
2 points: The question is entirely consistent with the character’s known decisions and the structure of the plot.
1 point: The question is generally consistent but has minor inconsistencies in detail.
0 points: The question is logically inconsistent with the character’s known decisions or the structure of the plot.
3. Challenge Level
Rule 3.1: Evaluators need to assess the plausibility of the incorrect options. Wrong options should be reasonably believable and attractive within the constraints of the character’s background and motivations, making the questions sufficiently challenging.
Scoring Guide:
2 points: All incorrect options are highly plausible, convincingly misleading.
1 point: Most incorrect options are reasonable, but one or two lack plausibility.
0 points: Incorrect options are obviously illogical and lack the ability to mislead.
4. Alignment with Character Motivation
Rule 4.1: Evaluators must assess whether the question correctly guides the testing model to step into the role and make a choice, i.e., testing if the model can replicate the real storyline’s choices. It is crucial that the character’s motivations, as articulated by literary experts, are a central component reflected in these questions.
Scoring Guide:
2 points: The question unambiguously points to a specific character decision point, accurately testing the model’s ability to role-play.
1 point: The question points to a character decision point to some extent, but the indicators are not clear enough, potentially reducing the accuracy of the model’s role-playing test.
0 points: The question fails to clearly define the character decision point, unable to effectively test the model’s role-playing ability.
Additional Notes:
1. Before starting the evaluation, each evaluator must understand the core motives and development axes of the character by reading summaries and analyses of the novels created by literary experts.
2. Ensure that evaluators are familiar with all background material before scoring any questions.
3. Evaluators should reference the analyses by literary experts of the characters to evaluate each of GPT-4’s multiple-choice questions, maintaining consistency of standards.
4. Application of the evaluation rules should be flexible and adapted to the specific context; scoring standards may be adjusted for special cases.
We evaluated the scores of each annotator and only retained the data with an average score of more than 6 points.
B.2 Manual Evaluation of Data Leak
Below are the manual evaluation standards for data leakage issues. We invited three annotators to evaluate the role stories continued by the model, with three levels:
Completely Consistent: The output is deemed "completely consistent" if the model restores the plot with high accuracy, which includes correct character interactions, significant plot points, and events as they happen in the book, despite the replacement of character names and other distinctive elements.
Partially Consistent: The output is "partially consistent" if the model produces content that partially matches the original plot, meaning it can still restore some elements of the story correctly, but there are noticeable discrepancies in the sequence of events, character behaviors, or outcomes not aligning with the text as it is in the book.
Entirely Different: The output is considered "entirely different" if the model’s content bears no resemblance to the original plot, character interactions, or events from the book after the entity replacement. This output suggests that the model is generating new content without relying on potential memorized data from its pre-training set.
The annotators need to select the appropriate level of story restoration for each character based on these three situations. The final proportion of each restoration level is shown in Table 5.
B.3 Manual Evaluation of GPT-4 Scores
In this section, we use GPT-4 to score whether the character profile generated by the model contains the manually annotated motivation. To ensure the quality of GPT-4 assessment, we sample 40 pieces of data and invite 3 native English-speaking college students for manual evaluation. We provide the annotators with the same scoring standards as GPT-4, and each annotator scores all data. We calculate the Pearson Correlation Coefficient Cohen et al. (2009) between the model and manual scores. The final results show that the Pearson Correlation Coefficient between GPT-4 and each annotator exceeds 0.7, and the correlation coefficient with the average score of annotators is 0.81. This indicates that GPT-4’s evaluation ability on this task can be comparable to that of humans.