### v2

# Character is Destiny: Can Role-Playing Language Agents Make Persona-Driven Decisions?

Abstract
Can Large Language Models (LLMs) simulate humans in making important decisions? Recent research has unveiled the potential of using LLMs to develop role-playing language agents (RPLAs), mimicking mainly the knowledge and tones of various characters. However, imitative decision-making necessitates a more nuanced understanding of personas. In this paper, we benchmark the ability of LLMs in persona-driven decision-making. Specifically, we investigate whether LLMs can predict characters’ decisions provided by the preceding stories in high-quality novels. Leveraging character analyses written by literary experts, we construct a dataset LifeChoice comprising 1,462 characters’ decision points from 388 books. Then, we conduct comprehensive experiments on LifeChoice, with various LLMs and RPLA methodologies. The results demonstrate that state-of-the-art LLMs exhibit promising capabilities in this task, yet substantial room for improvement remains. Hence, we further propose the CharMap method, which adopts persona-based memory retrieval and significantly advances RPLAs on this task, achieving 5.03% increase in accuracy. Resources are available at https://github.com/airaer1998/LifeChoice.
Figure 1: An example of LifeChoice. Given a character, a decision point and the preceding context, RPLAs are expected to reproduce the original decision. Typically, RPLAs are constructed by parsing the context into the character’s description and memory.
1 Introduction
The fault, dear Brutus, is not in our stars, but in ourselves, that we are underlings.
– Julius Caesar. Act 1, Scene 2.
With the recent advancements in large language models (LLMs) OpenAI (2023); Touvron et al. (2023), Role-Playing Language Agents (RPLAs) have emerged as a flourishing field of AI applications and research Chen et al. (2024). RPLAs are LLM-based AI systems that simulate assigned personas, reproducing their tones, knowledge, personalities and even decisions Park et al. (2023); Gao et al. (2024); Wang et al. (2024); Xie et al. (2024). They emulate various characters across extensive applications, including fictional characters in chatbots and video games Wang et al. (2023, 2024), as well as digital clones (Gao et al., 2023) or personalized assistants Xu et al. (2022); Salemi et al. (2024) for real-world individuals.
Can RPLAs reliably make decisions that align with their personas, as humans do? This question is vital for the practical usage of RPLAs, yet remains underexplored. Previous studies primarily investigate RPLAs’ character fidelity in terms of their tones (Wang et al., 2023) and knowledge (Shao et al., 2023), which could be readily replicated by existing RPLAs via style imitation and knowledge retrieval. However, these features are relatively superficial compared with the underlying thinking and mindset of characters. Recent efforts (Wang et al., 2024) study the personality fidelity of RPLAs, but they fail to capture the nuances and dynamics of characters’ mindsets. Hence, it remains an understudied question whether RPLAs could simulate persona-driven decisions, which challenges their comprehensive understanding of the personas and reasoning about unobserved behaviors.
In this paper, we systematically study the capability of RPLAs to simulate persona-driven decisions, based on characters from high-quality novels. In high-quality novels, characters’ life choices are carefully plotted and aligned with their personas. Hence, we introduce the LifeChoice dataset, which evaluate whether RPLAs can faithfully reproduce the characters’ life choices in the narratives. Specifically, LifeChoice comprises 1,462 character decisions from 388 novels, leveraging expert-written character analyses. Each sample is presented as a multiple-choice question with the preceding context before the decision point. As depicted in Figure 1, RPLAs are expected to identify and reason over relevant knowledge about the characters to simulate their decisions. The construction of LifeChoice primarily involves three steps: decision point selection, multiple-choice question construction, and manual examination.
Compared with previous methods for RPLA evaluation, our task and dataset benefit from higher-quality data and are more challenging. First, our questions and decisions are well-designed and closely aligned with the personas, since they are sourced from well-crafted narratives. Hence, our data establish solid ground truth for simulating characters’ persona-driven decisions. Second, our task is more challenging as it requires RPLAs to comprehensively understand and reason based on the personas, including their knowledge, experiences, and personalities. Specifically, LifeChoice poses the following challenges: 1) Long-context understanding, where RPLAs need to identify sparse relevant motivations from massive character contexts. 2) Temporal intelligence, where RPLAs should intelligently adapt to the dynamic evolution of characters and environments. 3) Intricate motives, where RPLAs are required to reason through complex and entangled backgrounds and motives to arrive at the decisions.
We conduct extensive experiments to evaluate RPLAs on LifeChoice. Our experiments cover various LLMs and different RPLA frameworks, including memory-enhanced agents, long-context LLMs, and our proposed method CharMap towards better simulation of persona-driven decisions. The results demonstrate that existing RPLAs have shown a promising accuracy of up to 62.92% on LifeChoice. Furthermore, CharMap significantly enhances RPLAs on this task, achieving an accuracy of 67.95%, which exceeds previous baselines by 5.03%. However, compared to the human performance of 92.01%, there is still significant room for improvement. Meanwhile, we observe that both well-summarized character descriptions and accurate memory retrieval are crucial for RPLAs.
In summary, our contributions include:
  • We propose to explore RPLAs’ ability in simulating persona-driven decisions, which is crucial for future RPLA applications and challenges existing RPLAs.
  • We delicately craft LifeChoice, the first benchmark for persona-driven decisions of RPLAs, based on characters’ life choices from high-quality novels. Besides, we propose CharMap, which adopts persona-based memory retrieval for better decision-making of RPLAs.
  • Based on LifeChoice, we conduct extensive experiments. The results demonstrate the promising performance of RPLAs in decision simulation. Then, we analyze and compare methodologies for RPLA development, and show the effectiveness of CharMap.
2 Related Work
Character Role-Playing
Early research on character-related studies focuses on character understanding. Brahman et al. (2021) attempts to predict a specific character through the text of the novel. Yu et al. (2022) provides dialogues from movie scripts for the model to examine and then asks it to identify the character who speaks each passage. With the enhancement of model abilities, some work attempts to make the model simulate complex role-playing. Li et al. (2023) analyzes 32 anime characters using 54k dialogues and personality traits. They use sentence embeddings for dialogue selection and evaluation. Zhou et al. (2023) uses identity, interests, and relationships, collecting AI behaviors for imitation and using character data for fine-tuning. They evaluate model consistency and linguistic style. Wang et al. (2023) creates a dataset for script characters and evaluates role-playing quality based on speaking style imitation and role-specific knowledge. These studies make a chatbot for a certain character, but they focus more on imitating the character from the perspective of dialogue, which is a shallow imitation. Our study aims to evaluate RPLAs from the perspective of behavior and decision-making, which further challenges the model’s understanding of the roles.
Personal LLM assistants
With the rapid development of artificial intelligence technology, there are now many personal intelligent agents embedded in mobile devices, providing personalized services through analyzing user data and equipment Kaplan and Haenlein (2019); Hoy (2018). These agents can model the user’s profile and preferences through the user’s historical data Gurrin et al. (2014); Dodge and Kitchin (2007), such as extracting personality from the user’s record text Majumder et al. (2017); Štajner and Yenikent (2020), reading emotions from the user’s image data Jaiswal et al. (2020); Zad et al. (2021), modeling preferences from historical interaction information Tang et al. (2019); Li et al. (2018), and pushing notifications from smart phones Li et al. (2018). These memories can enhance the model’s decision-making and reasoning, bringing a better personal experience for users. However, obtaining real user memory data is difficult and violates privacy. We model characters from historical data in high-quality novel texts, allowing the model to restore the real choices in the storyline based on the previous text, providing the first benchmark for the wide testing of personal intelligent agents.
3 Dataset and Task Setups
3.1 Dataset Construction
Book: Les Misérables
Character: Jean Valjean
Context:
In 1815 Monsieur Charles-François-Bienvenu Myriel was Bishop of Digne. He was then……Jean Valjean reflections gave him a sort of frightening aspect. He was subject to one of those violent inner tearings, which was not unknown to him.
Scenario:
In the courtroom, an innocent man was wrongfully accused, because he bore a resemblance to Jean Valjean. If Jean Valjean did not come forward, this innocent man would be sent to the gallows in his place. At this time, Jean Valjean had transformed his identity and become a respected town mayor, and he had also adopted a young girl named Cosette, with whom he had a new life.
Question:
You will play the role of Jean Valjean. What will you choose to do when you discover that man is about to be convicted due to being mistaken for you?
Options:
A. Keep silent, letting an innocent person take the punishment in one’s place.
B. Persuade the person to run away, in order to protect both from the disaster of jail.
C. Go to court and reveal the truth, sacrificing oneself to save the innocent person.
D. Look for legal loopholes, trying to save both the person and oneself.
Correct Answer: C
Motivation:
[Values and Beliefs] Jean Valjean is a person who values honesty and justice, possessing a strong sense of morality and righteousness. He decides to turn himself in to save another innocent person, fulfilling his inner need for morality and justice.
Table 1: Case study of LifeChoice. A complete set of data includes book, character, scenario, question, options, correct answer, motivation, and input.
Dataset Source Context Length Task Format Has Explanation
TVSHOWGUESS TV show transcripts ~50k Character Identification
ROCStories Commonsense short stories ~100 Character Behavior Prediction
LiSCU Literature ~1000 Character Identification
LifeChoice Literature ~150k Character Behavior Prediction
Table 2: Comparison between LifeChoice and previous character understanding benchmarks: data source, context length, task format, and whether the benchmark has explanations.
We construct a comprehensive dataset called LifeChoice. As shown in Table 1, the sample for each decision point includes the preceding context $p$ from the original book, the current scenario $s$, a question $q$ outlining a decision faced by that character $c$, a list of options $a=\{a_{i}\}_{i=1}^{4}$, the correct answer $y$, and the motivation $m$ explaining the character’s choice. Our data is sourced from the website Supersummary [^1] https://www.supersummary.com/, which provides three pieces of content written by literary experts: key character descriptions, chapter summaries, and book analyses. We contact the website and obtain authorization to use the data for academic research. The dataset construction comprises the following three main steps:
Selecting Decision Points
To prevent data leakage, we first filter novels on the site using the following criteria: (1) The narrative must exclude non-fiction genres like biographies or documentary literature. (2) The narrative perspective must be in the first or third person. (3) The progression of narrative time should be linear, avoiding stories with complex timelines or flashbacks. (4) Exclude overly popular books, as measured by a high number of reviews on literary review websites. For each book that passes these filters, we provide GPT-4 with content written by literary experts, requesting it to output each key character’s life choice decision points and their corresponding gold motivations. Additionally, we also ask GPT-4 to output the chapter numbers related to the decision based on the extracted motivations. As shown in the example in Figure 1, the literary expert’s analysis of the book suggests that Michael Corleone’s motivation for choosing to assassinate the enemy includes both avenging his father and witnessing the collusion between the police and the enemy, which exposes him to the darker side of the government. We then identify two corresponding chapters in the original book based on these motivations, providing more refined data for constructing multiple-choice questions.
Constructing Multiple-Choice Questions
We input the content written by literary experts and the corresponding chapters identified based on motivation into GPT-4. Our goal is to generate multiple-choice questions that capture the complexity of the characters’ decision-making processes. The correct option reflects the decision made by the characters in the original books, whereas the distractors are designed to be plausible for an arbitrary person. As shown in the example in Figure 1, Michael Corleone can ask for help from the government because he was once a Navy officer who trusted the government. However, in the preceding text, Michael witnesses the dark side of the government, so he ultimately chooses to stab the police.
Manual Examination
We invite ten native English-speaking university students to filter the data and pay them according to local minimum wage standards. We supply the annotators with content written by literary experts and the multiple-choice questions, asking them to assess whether the model-created questions are challenging and reasonable. They are also tasked with filtering out data they deem low quality.
Ultimately, we collect 1,462 characters from 388 books and their corresponding life choices. Table 1 shows a complete data example. The specific prompts and more detailed data construction process can be found in Appendix A.
Figure 2: Statistics of motivation types in LifeChoice, with the first words for each motivation type.
3.2 Dataset Analysis
We refer to the drama theory of Aristophanes Sommerstein (2013); Silk (2002) as the system prompt and use GPT-4 to classify the motivations for character decisions into two meta-motivations and several accompanying sub-motivations:
Character-driven Motivation
Character-driven behavior revolves around the character’s inner world, personality, and transformation. Sub-motivations of character-driven behavior include Personality and Traits, Emotions and Psychological State, Social Relationships, Values and Beliefs, and Desires and Goals.
Plot-driven Motivation
Plot-driven behavior stems from a series of external events and conflicts unfolding. Characters often react passively within a larger narrative structure, with their actions led by external events. Sub-motivations of plot-driven behavior include External Conflicts, Tasks and Goals, Puzzles and Secrets, Pursuits and Escapes, Exploration and Discovery, Power and Control, and Intrigue and Betrayal.
Note that each topic is assigned one category of motivation. Figure 2 shows the proportion of different motivations. Detailed introductions for each sub-motivation are in Appendix B.2.
3.3 Task Setups
This task can be formulated as $P(y|x)$. Given the input $x=(p,s,c,q,a)$, the RPLA needs to identify the correct choice $y$ that aligns with the character’s decision in the narrative. For evaluation, we directly use the accuracy of multiple-choice question answering. As shown in Table 2, compared to other character understanding tasks, LifeChoice requires understanding the character through a more extended context to make decisions. RPLAs must locate relevant information related to the current scene in vast personal data. This behavior demands a more profound understanding of the characters.
4 Experiments
Profile Construction Role-Playing Model ACC +motivation
Description Construction
Hierarchical merging LLaMA-3 42.10 83.09
GPT-3.5 39.85 80.00
GPT-4 45.43 85.24
[…]
GPT-3.5 41.06 81.63
GPT-4 47.02 86.47
[…]
GPT-3.5 52.04 86.33
GPT-4 55.17 90.23
Memory Retrieval
BM25 GPT-4 26.08 75.88
Embedding GPT-4 35.66 78.24
Description & Memory
Direct concatenation LLaMA-3 57.02 92.04
Mixtral 58.56 91.75
Claude-3.5 62.85 96.45
Gemini-1.5-pro 57.16 91.38
GPT-3.5 55.62 90.39
GPT-4 62.92 95.46
[…]
Mixtral 65.02 92.05
Claude-3.5 67.13 96.90
Gemini-1.5-pro 63.94 91.39
GPT-3.5 61.62 90.95
GPT-4 67.95 96.87
Table 3: Results of different LLMs on LifeChoice. ACC refers to the decision accuracy. +motivation refers to the results after providing the motivations behind character decisions, which are extracted from expert analyses by GPT-4.
Because our inputs generally exceed 100k, it is difficult for LLMs to handle them directly. Therefore, our approach is divided into two steps: 1) Character Profile Construction, which includes the character’s description and memories; 2) Reasoning for Decisions, where different LLMs use the constructed profile to answer the questions.
4.1 Character Profile Construction
As shown in Figure 1, the character profile consists of two parts. The first part is the character’s description, including their personality, experiences, hobbies, etc. The second part is the character’s memories, specific segments from the preceding text. Below, I will detail the methods for constructing these two parts:
Description Construction
We adopt two automatic methods to construct character descriptions: (1) Hierarchical merging Wu et al. (2021): Books are divided into chunks that fit within the LLM context window. The LLM summarizes each chunk, then merges and summarizes adjacent summarized chunks iteratively to produce the final description. (2) Incremental updating Chang et al. (2023): Books are divided into chunks and summarized sequentially, and the description is updated and refined incrementally by concatenating summarized chunks. The summarization model for both automated methods is GPT-3.5. Additionally, using the (3) expert-written descriptions from Supersummary, we employ GPT-4 to identify the positions of the decision points and truncate the text, providing only the data before these points. All descriptions are kept within 5k tokens, the maximum for human-written descriptions.
Memory Retrieval
We use two memory retrieval methods: (1) BM25 Robertson et al. (2009): Scores documents based on term relevance and length, optimizing retrieval using term frequency and distribution. (2) Embedding-based retrieval: Uses dense vectors representing documents and queries to assess semantic similarity through vector distance. For the embedding model, we use OpenAI’s text-embedding-ada-002Neelakantan et al. (2022) model.
Figure 3: An overview of CharMap, a two-step scenario-specific character profile building approach.
Description & Memory
Using only Description or Memory alone may lead to information loss (Wang et al., 2024). Therefore, we also experiment by combining the results of both methods to form the character’s profile. We adopt two methods: (1) Direct concatenation: This method concatenates the results from both approaches by prompting the model to role-play the corresponding character. By default, it uses the results from Human Description and Embedding retrieval. (2) CharMap: To better utilize the information in the Description, we propose CHARacter MAPping Profile Synthesis (CharMap), constructing a more scenario-specific profile in two steps. As shown in Figure 3, first, after obtaining the description, we input it along with the question into the model, asking it to locate the plot in the Description relevant to the current scene based on the question. Second, we use these episodes as queries to retrieve related memories and then input them into the inference model and the description. This leverages the overall character storyline in the description, thereby better retrieving related memories.
4.2 Reasoning for Decisions
After compressing the original input $x$ into a character profile, we feed it into the LLMs. For methods using only description or memory, we use GPT-3.5, GPT-4, and LLaMA-3Team (2024b). For methods using both, we also include Claude-3Anthropic (2024), GeminiTeam (2024a), and Mixtral Jiang et al. (2024). For all these models, we adopt the official instruction formats where available  [^2] The versions in this paper are gpt-3.5-turbo-1106, gpt-4-1106-preview, Llama-3-70B-Instruct, Claude -3.5-Sonnet, Gemini-1.5-pro and Mixtral-8x7B-v0.1 respectively..
5 Analysis
Raw text Concat. CharMap
GPT-4 - 63.05 68.10
human 92.01 66.82 74.78
Table 4: Results of the human evaluation. Concat. refers to the direct concatenation of Description and Memory.
In the experiments, we wish to answer two research questions: RQ1) Can LLMs make decisions based on historical data? RQ2) What influences the decision-making of LLMs?
5.1 Can LLMs Make Decisions Based on Historical Data?
Analysis of Model Results
Table 3 presents the accuracy results of different RPLA methods on the LifeChoice. Additionally, we evaluate the results when the model is provided with gold motivation, and several observations can be made: First, the method that uses both Description and Memory surpasses the one that uses only one, suggesting that both holistic and detailed data of key characters are essential in final decision-making. Second, when gold motivation is provided, the accuracy consistently exceeds 80%, indicating the rationality of these motivations in the data. Third, the performance gap among different LLMs is not significant while reasoning the answer. This indicates that the main factor for the result is the generated profile rather than reasoning ability. Last, CharMap outperforms the method that directly concatenates Description and Memory by 5.03%, proving its effectiveness. This scenario-specific profile better assists RPLA in decision-making.
Humans are Good Decision-makers
We invite three native English-speaking university students to take a test in which we select six novels they have never heard of before. Each novel has between 3 to 5 characters and their corresponding multiple-choice questions. We provide each person with three data sets for each key character in two books: the full original text before the decision point, direct concatenation Description and Memory result, and the result from CharMap. As shown in Table 4, compared to direct concatenation, the CharMap results are easier for humans to understand. Additionally, humans slightly outperform GPT-4 in reasoning answers based on the profiles, indicating that humans can understand subtle character decisions better than models. When given the raw text, humans can achieve an accuracy rate of 92.01%, suggesting there is still significant room for improvement in RPLA methods.
Figure 4: The impact of the number of book reviews on accuracy in LifeChoice, with new books being those not present in the training corpus of LLMs.
Mitigation and Analysis of Data Leakage
Data leakage is a significant challenge since our data might appear in the model’s pre-training corpus. During the data collection phases in section 3.1, we adopt various preventive measures. For evaluation, we employ an entity replacement strategy, substituting character names, locations, and other entities with placeholders. We believe data leakage relates to the amount of relevant corpus used during LLM pre-training, with more popular books having more related corpus. To verify this, we use the number of reviews on the book review website [^3] https://www.douban.com/ to indicate a book’s popularity and evaluate the results of books with different review counts on LifeChoice. We use CharMap to build profiles and GPT-4 as the role-playing model, sampling thirty books with different numbers of reviews, including thirty books not in the LLMs’ corpus (published after November 6 for gpt-4-1106-preview). As shown in Figure 4, the model’s accuracy significantly improves when the number of reviews exceeds 5,000. In contrast, books with fewer than 5,000 reviews show slight fluctuation and results similar to those not in the LLMs’ corpus. Therefore, it can be considered that for books with a low number of reviews, data leakage has little impact on CharMap. In section 3.1, we use 5,000 reviews as a threshold to filter the books.
Analysis of Long-Context LLMs
Long context is an essential feature of LifeChoice, and directly using long-context models for role-playing is an exciting topic. Making decisions based on extensive context tests a model’s ability to understand global data and reason from a character’s perspective. We evaluate two long-context models: Claude3-sonnect and kimi-chat. As shown in Table LABEL:tab:longcontext, although the performance of long-context models is not as strong as CharMap, they still demonstrate potential in role-playing. LifeChoice, as a task requiring multiple reasoning points and an overall understanding of the context, can also serve as a vital benchmark for evaluating long-context models.
5.2 What Influences the Decision-making of LLMs?
Figure 5: Heatmap of the impact of motivation types on the results. The results are predicted from the Incremental updating, the embedding-retrieved memory, the direct concatenation of both, and CharMap. The role-playing model uses GPT-4.
The Impact of Motivation Types
In line with the motivation types presented in Section 3.2, we examine how different types of motivation influence characters’ decision-making. For profiles, we evaluate four methods: the Incremental updating, the embedding-retrieved memory, the direct concatenation of both, and CharMap. For reasoning, we use GPT-4 uniformly. The results are shown in Figure 5. We find that tasks requiring coherent reasoning, such as puzzles and mysteries, are not well answered for all methods. This might be because these questions need multi-step reasoning and details from various memories. Moreover, plot-driven questions have lower accuracy when descriptions are used only for the profile. Conversely, character-driven questions are challenging to answer when relying only on memories. We believe this is because character summaries in descriptions better capture the overall essence of the characters, while memories provide direct access to relevant events.
Figure 6: The result of the impact of different novel genres on accuracy.
The Impact of Novel Genres
We use the genre tags from novels on the website to analyze the accuracy of character selection across different genres. We conduct experiments on the the direct concatenation of description and memories, and the role-playing model using GPT-4. As depicted in Figure 6, the accuracy of science fiction, fantasy novels, and romance novels is quite high. This could be because the characters in these novels are often stylized or have fixed creative patterns and archetypes. In contrast, crime and mystery novels perform poorly, which might be because they involve complex logical chains, and characters in these novels frequently take abnormal actions. Further details about each genre and the complete table can be found in Appendix B.1.
Figure 7: Analysis of whether character selection will change. The x-axis represents the input length relative to the point truncation.
The Impact of Temporal Data
If faced with the decisions of years past at this moment, would you make the same choices? We conduct a study on this matter. Specifically, we randomly sample 40 characters, half character-driven, and half plot-driven. We split the content preceding the decision points into five equal sections and used these various content lengths as input. We conduct experiments on the combination of human description + embedding-retrieved memories, and the role-playing model is GPT-4. As shown in Figure 7, in the early stages, the accuracy of most characters’ decisions is close to random (25%), potentially due to insufficient information. As more information becomes available, the characters’ decisions tend to be closer to the correct choice. For character-driven decisions, accuracy tends to be stable. For plot-driven, the accuracy rate may change abruptly. This could be due to the relatively stable characteristics of a character, while some sudden events may greatly influence the final choices of the character.
6 Conclusion
In this work, we propose the first task to evaluate the decision-making of RPLAs, testing whether LLMs can accurately reconstruct decisions using historical data. We construct LifeChoice, which includes 1,462 characters from 388 books and their life choices. Extensive experiments on LifeChoice demonstrate the promising performance of RPLAs in decision simulation. Additionally, we propose CharMap, which uses persona-based memory retrieval to enhance decision-making. We hope this work provides better evaluation benchmarks for RPLAs and directs the future development of personal LLM assistants.
Limitations
In this paper, we primarily investigate whether fictional characters can recreate their choices within a book. Although we have controlled the quality of the novels, there may still be issues with the plot and characters since the author designed the storyline, which can result in illogical choices within the book. Furthermore, as our research focuses mainly on fictional characters, there is a certain gap compared to real-world humans. For example, the author fictionalizes some story backgrounds, which may impact the model’s generation results.
Ethics Statement
Use of Human Annotations
Our institution recruits annotators to implement the annotations of motivation recognition dataset construction. We ensure the privacy rights of the annotators are respected during the annotation process. The annotators receive compensation exceeding the local minimum wage and have consented to use motivation recognition data they process for research purposes. Appendix C provides further details on the annotations.
Risks
The LifeChoice dataset is sourced from novels and analysis data written by human literary experts. However, we cannot guarantee that this data is entirely free from toxic or discriminatory language. Additionally, our data is generated by GPT-4, which may introduce some inherent biases and hallucinations of the model.
Appendix A Prompts
In this section, we provide the key prompts we used, including prompts for selecting decision points, locating the node’s position, constructing multiple-choice questions, system prompts for role-playing characters, and the prompt of CharMap.
A.1 Selecting Decision Points
As mentioned in section 3.1, the first step in constructing our data is selecting the character’s Decision Points. In this step, our input data consists of all raw data from Supersummary, including the current character’s description, all chapter summaries, and book analysis. We provide this data to GPT-4, requesting it to output three types of data: the character’s decision point, the corresponding gold motivation, and the chapters related to this choice. Table 5 presents the character’s description, Table 6 presents the chapter summaries, and Table 7 presents the book analysis, these examples are all from the 2024 novel "A Calamity of Souls.". Table 8 illustrates the prompt for selecting decision points.
A.2 Locating the Node
Furthermore, for the character’s decision points, we provide the previously identified decision points and their corresponding chapters from the original book to GPT-4. This allows it to precisely determine the position in the original book that should be segmented, helping to avoid data leakage. Table 9 shows the prompt for locating.
A.3 Constructing Multiple-Choice Questions
After selecting the Decision Points, our next step is constructing Multiple-Choice Questions. In this step, our input data consists of all the input and output data from the previous step(Appendix A.1). We ask GPT-4 to construct a multiple-choice question regarding the character’s decision based on this data, outputting the scenario in which the character is situated, the question, and four options. Table 10 shows the prompt for constructing multiple-choice questions.
A.4 System Prompts for Role-Playing
The prompt for role-playing as the character can be found in Table 11.
A.5 Prompts for CharMap
The prompt for CharMap can be found in Table 12.
Appendix B Dateset Details
B.1 Categories of novel
Below is a complete classification of novel genres, from the literary experts at the Supersummary website:
Mystery Novels: The mystery genre includes general mystery, noir mystery, historical mystery, police procedural mystery, and supernatural mystery.
Thriller Novels: The thriller genre includes supernatural thrillers, historical thrillers, environmental thrillers, medical thrillers, legal thrillers, political thrillers, military thrillers, and espionage stories.
Science Fiction Novels: Science fiction stories take place in the future or the past but are almost always set in a dimension different from our present. They are characterized by entirely new, imagined realities and universes, where the setting is indispensable. High technology also plays an important role in these stories. Space opera, romantic science fiction, military science fiction, alternate history, dystopian and utopian tales, as well as steampunk, are considered sub-genres of science fiction.
Romance Novels: Romance novels feature romantic relationships between at least two people, characterized by tension and desire. Romance novel themes include supernatural romance, contemporary romance, historical romance, western romance, gothic romance, regency romance, and romantic suspense.
Fantasy Novels: Fantasy stories are centered around mythical kingdoms and magic. Fantasy novel genres include contemporary fantasy, traditional fantasy, horror fantasy, weird fantasy, epic fantasy, historical fantasy, dark fantasy, urban fantasy, and anime fantasy.
Action Adventure Novels: Action-adventure novels place the protagonist in various realistic dangers. This is a fast-paced genre where the climax should provide some form of thrill for the audience or reader.
Speculative Novels: Speculative fiction is characterized by overlapping with our world but differing in key aspects, introducing "what if" scenarios.
Mystery Thriller Novels: Mystery thriller stories are usually filled with suspense, with one or more characters’ lives in danger. In gripping scenes, these characters are often chased and manage to escape narrowly.
Young Adult Novels: Young Adult fiction, commonly abbreviated as YA, is intended for teenagers aged 12-18. Most YA novels feature coming-of-age stories, often with elements of science fiction or fantasy.
New Adult Novels: New Adult novels target college-aged adults and usually explore stories of first adventures on one’s own.
Horror and Supernatural Novels: Horror, supernatural, and ghost story genres aim to scare the reader and audience by playing on common fears. The protagonist usually has to overcome supernatural threats, and the stories often include supernatural elements.
Crime Mystery Novels: Crime mystery stories focus on a central problem or crime to be solved, or a mysterious event that must be answered. Throughout the story, the reader or audience and characters are given clues that help the protagonist eventually find the solution.
Detective Novels: In detective fiction, a common element is a police officer or detective embarking on solving a crime. The plot is filled with evidence gathering, forensic studies, and legal drama.
Historical Novels: Historical novels are fictional stories set against the backdrop of real historical events or historical settings. Historical fiction may also portray real historical figures.
Western Novels: Stories with a western theme take place in the old times of the American West, filled with adventure, cowboys, and pioneers. There are also Italian western novels, Asian western novels, space westerns, and other stories about the American West.
Family Saga Novels: Family saga novels typically tell the stories of several generations of family members dealing with family affairs, family curses, and family adventures. These stories usually follow a timeline and deal with conflicts in the present.
Women’s Novels: Women’s fiction plotlines revolve around the challenges and crises that women face in real life, including interpersonal relationships, work, family, politics, and religion.
Magical Realism Novels: Magical realism stories take place in the real world but have characters who take magical elements for granted. These mystical elements do not exist in real life, but they are perfectly normal in magical realism.
B.2 Categories of motivations
Below are the motivations for each topic and their corresponding proportions:
Character-driven motivation
Character-driven narrative is centered on the inner world, growth, and transformation of characters. In character-driven stories, the progression of the plot and the resolution of conflicts are often propelled by the characters’ personalities, desires, fears, and psychological development. Such stories typically delve deeply into the characters’ mental states and development, focusing on how characters influence each other and how their actions reflect their inner emotions and thoughts. The choices and changes of the characters serve as the main engine for the story’s development, influencing the direction of the plot. Sub-motivations of character-driven behavior include:
Personality and Traits: (27.12%) These refer to a character’s characteristics such as being introverted, extroverted, brave, or guilt-ridden, which influence their choices and lifestyle.
Emotions and Psychological State: (7.53%) A character’s emotional responses, psychological traumas, or sense of personal well-being are key elements that drive the story forward.
Social Relationships: (6.31%) The character’s status and changes in family, love, friendship, or other social connections can propel the story’s development.
Values and Beliefs: (27.12%) The character’s moral convictions, religious beliefs, or life philosophy can serve as motivation for action.
Desires and Goals: (7.22%) Personal desires, career aspirations, or specific life goals of a character are pivotal in advancing the plot.
Plot-driven motivation
Plot-driven narrative emphasizes the creation and resolution of external conflicts in the story. In such stories, the driving force of the plot comes from a series of events and conflicts themselves, while characters are often the responders to these events. Plot-driven stories typically highlight tense drama, complex plot structure, and frequent changes in external actions, rather than changes in the character’s internal world. In this type of narrative, characters may act in response to the demands of the plot, rather than the plot following the development of the characters’ inner world. Sub-motivations of plot-driven behavior include:
External Conflicts: (8.76%) Conflicts from the outside world, such as war, natural disasters, or social upheaval, can propel the plot.
Tasks and Goals: (4.7%) Tasks or specific goals that characters must accomplish often become the driving force behind the story’s progression.
Puzzles and Secrets: (7.22%) Secrets that need revealing or mysteries that need solving can form the core of a story.
Pursuits and Escapes: (4.25%) Characters might chase something (e.g., power, wealth, knowledge) while avoiding or fleeing from certain situations (e.g., pursuit, personal past).
Exploration and Discovery: (3.66%) Characters’ adventures or discoveries in new realms (physical, scientific, or spiritual) can move the plot forward.
Power and Control: (4.81%) The pursuit or struggle for power and control often serves as motivation for characters.
Intrigue and Betrayal: (4.09%) Complex plots and betrayals can catalyze the progression of the story.
Appendix C Manual Annotation
This is a supplement to Section 3.1. After constructing the multiple-choice question data using GPT-4, we perform a manual examination. For each annotator, we provide key character descriptions, chapter summaries, and book analyses written by human literature experts on the Supersummary. Each annotator is asked to score the questions constructed by GPT-4 based on the rules shown in Table 13. We evaluated the scores of each annotator and only retained the data with an average score of more than 6 points.
We provide compensation based on the local minimum hourly wage for all individuals involved in the annotation.
Appendix D Future direction
Building personal agents for everyone is an exciting topic. We have explored how fictional characters can determine their subsequent actions based on historical data, proposing possibilities for combining role-playing with personalized models. We believe there are additional directions to explore:
  • Real-life version of LifeChoice The behavior of fictional characters often stems from the author’s design, which can lead to logical inconsistencies. In contrast, real-world human behavior data do not have this issue. How to construct real human historical data and related behavior data is a question worth exploring.
  • Improving RPLA performance in life choices Although CharMap has achieved decent results, better methods are needed to balance reasoning efficiency (which depends on the length of input tokens) while achieving superior outcomes.
  • More complex downstream decision tasks The decisions we select are often significant choices for fictional characters, resulting in a large decision space without a fixed task framework. Identifying more systematic tasks by integrating social sciences is a challenge that needs to be addressed in the future.